Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
karathanasis, paoletti, menendez, fiori, tringali - 17:51 Friday 25 October 2019 (47394) Print this report
Task ENV 06: Magnetic coupling at SDB1
Two days ago, during our ENV shift, we performed close-field magnetic injections on CEB.
First we successfully reproduced the results from #46228 for the 368 Hz line on detection.

Then, we went up to the DET terrace and we performed several line injections at frequency
368 Hz using the small coil on the racks and cables.
We noticed an effect on DARM when the coil was located near some electronics,
like the Box LVDT and quantrants.(see fig. 1)

Moreover, we tried to check if the coupling was due to the cables or the electronics.
To do this, we separated the STP cables behind the rack and injected the line, moving the coil
very close to each one of them, while keeping the others further away. (see fig. 2)

In addition, we decided to inject white magnetic noise and noticed an effect on wide range of
frequencies.(see fig. 3)

Then, we moved to the EE room and performed white noise injection nearby SIB1. No effect
was found. For this reason, we decided to increase the level of the injection and
found some structures around 150 Hz.(see fig. 4)
Additionally, we injected the 368 Hz line to the electronics in the room and no effect was
found.

Lastly, we decided to move the small coil to the south flange of the DET tower and injected
the same white noise as before.(see fig .5)
Images attached to this report
Comments to this report:
karathanasis - 12:00 Monday 28 October 2019 (47431) Print this report
Attach you can find the log file of the shift.
Non-image files attached to this comment
mwas - 10:12 Tuesday 28 July 2020 (49110) Print this report

Looking back at the magnetic injections done around SDB1 in October 2019.

Figure 1, shows in blue a calm time, in purple with the magnetic broad band injections next to the SDB1 quadrant and LVDT electronics and in brown next to the south flange of the tower.

The purple curve shows that the magnetic noise injection in the electronics rack increases the noise in the quadrant error signal by 2 orders of magnitude, and in the feed back signal by 1 order of magnitude. But this is still 2 orders of magnitude smaller than the DAC noise of these actuators (see figure 2 taken from VIR-0369A-20). So the coupling path is not the feedback being spoiled.

Figure 3 (from VIR-0369A-20) shows that there are mechanical resonances at 159Hz and 167Hz that can be excited when shaking the bench with the local controls. So the suggestion given in VIR-1029A-19 that magnetic fields are shaking the SDB1 bench by acting directcly on the bench magnets seems plausible. Although it would be important to have a noise projection from the ambient magnetic noise, to see if the amplitude of the noise projection is sufficiently close to the noise curve. The 167Hz peak is actually visible in the noise curve, which cannot be explained by the local control DAC noise, but maybe could be explained by the ambient magnetic noise.

A mitigation could be to replace the magnets on SDB1, and reallocated some of the correction signal to the magnets on the SDB1 marionette. However, first the transfer function between the marionettes actuator and the bench needs to be measured, to see if there is any filtering provided in the vertical degree of freedom by the 3 wire connecting SDB1 to the marionette, which do not provide any vertical angular isolation at the pendulum mode, but might provide some isolation at higher frequencies. This measurement could be done as soon as the SDB1 electronic racks are repositioned and re-connected.

 

Images attached to this comment
bulten - 12:14 Wednesday 29 July 2020 (49116) Print this report
SBE_SDB2 position control turned red since the actuator is now above 6 V; I adjust the stepper motor 4 for the LVDT-1 vertical keystone to get the controls in range again. I will start this at 10:15 UTC.
This operation will be visible in the trend plots. During the first day, the actuator voltage changed by 6 V and the LVDT positions of filter-0 and filter-1 y by about 600 micrometers, see attached trend plots.
Non-image files attached to this comment
bulten - 13:23 Wednesday 29 July 2020 (49117) Print this report
I adjusted the keystone positions of SDB2 filter-1 and filter-0 with the stepper motors from UTC 10:20 to 10:45; the SDB2_LC_Y table had some oscillations during that period.
The vertical actuator on filter-0 went from -6V to -1.8V, I think that that will be enough to cope with another day of thermal transient study if necessary. Since the LVDT-y position of SDB2 did not start at 0, but rather at +0.6 mm, it may be that we do not have to adjust the fishing rod for filter-1 when the powersupply 8 on the bench is switched on again, but
we do need to re-adjust the fishing rod for filter-0 since I expect that the actuator will go back to +4V when the powersupply is switched on again.
The stepper motor drivers were on from UTC 10:20 till UTC 10:46, I adjusted motor 4 for about 18 minutes and motor 3 for about 2 minutes after that. For motor-4, the relative LVDT-positions change (filter-1 is pushed down) but for motor-3 only the actuation voltage SBE_SDB2_ACT3_raw changes, since the vertical position of filter-0 is controlled and the fishing rod just changes the amount of force needed. I add the plots of the thermal transients from the moment that the DAQ power was switched off (July 28, UTC 6:43) till a few minutes after the stepper motor activations, and a zoom for the hour I adjusted the stepper motors.
Images attached to this comment
Search Help
×

Warning

×